b. Please compute cost per unit of each product on the new method of allocation

Writers' Company produces 2 products presently

The information related to manufacturing is given below.

WRITERS' Company

Yr 2018-19

Fountain Pen

Ball point

Units

10000

20000

Area occupied (sq feet)

7000

8000

Variable cost – Direct

32

25

Fixed Cost - Direct (per unit)

10

8

Rent - Fixed Cost per month (Indirect

- common for all products)

3,00,000.00

1
Expert's answer
2020-12-23T07:17:47-0500

b) The cost per unit is shown in the table below,


"Working (a) \\\frac{(300000\times7)}{15}=140000\\
\frac{140000}{10000}=14\\
Working(b)\\
\frac{(300000\times8)}{15}=160000\\
\frac{160000}{20000}=8"